# https://uni-fl.github.io/

### v1

@@ 5.3 Ablation Study
Figure 6: Illustration of PeFL’s impact on structure optimization. The instance segmentation model(SOLO) provides more effective feedback on structure generation issues such as distorted limbs, missing key objects, and subject ambiguity during fine-tuning.
Figure 7: The ablation study on the PeFL optimization. We compared the generated results utilizing diffusion loss and PeFL with the same perceptual dataset. (a) The effect of PeFL on style optimization. (b) The effect of PeFL on the structure optimization.
How PeFL works.
To gain a better understanding of how PeFL works, we take the example of structure optimization with PeFL and visualize the intermediate results. As shown in Fig.6, the instance segmentation model effectively captures the overall structure of the generated object and successfully identifies the structural defects, such as the distorted limbs of the little girl, the missing skateboard, and the surfboard, and the ambiguity elephant and horse. Instead of assigning equal importance to each pixel with naive diffusion loss, this type of feedback enables the diffusion model to focus more on the specific structured concepts. We showcase some generation results after optimization style and structure via PeFL in Fig.7. It is evident that the PeFL significantly boosts style generation(e.g. ’frescos’, ’impasto’ style) and object structure optimization(e.g. the woman’s glasses, ballet dancer’s legs) compared with the naive diffusion loss.
Effect of decoupled feedback learning.
To verify the importance of this decoupled aesthetic strategy, we conduct experiments by fine-tuning the SD1.5 model using a global aesthetic reward model trained with all the collected aesthetic preference data of different dimensions. As depicted in Fig.8(a), due to alleviate the challenge of abstract aesthetic learning, the utilization of decoupled aesthetic reward tuning resulted in generation results that were preferred by more individuals, surpassing the non-decoupled way by 17%. Fig.8(a) also shows that the active prompt selection obtained a higher preference rate(54.6% vs 45.4%), which demonstrates the importance of the prompt selection strategy.
Figure 8: (a) Human preference evaluation of different components in UniFL with SD1.5. (b) Human preference evaluation of the generated video via plugging the AnimiateDiff module into SD1.5(light green) and SD1.5 optimized by UniFL(dark green).
Figure 9: Experiment of SD1.5 without and with the adversarial objective. (a) The intermediate results in 20 steps inference. (b) The 4 steps inference results.
@@
How adversarial feedback learning accelerates. UniFL introduces adversarial feedback learning for acceleration, and the acceleration results even exceed the non-acceleration model in some scenarios; according to our experimental observations, the reason for the acceleration and the significant performance can be attributed to two potential factors: (i) Adversarial training enables the reward model to continuously provide guidance: As shown in Fig.9(a), the diffusion models with traditional feedback fine-tuning often suffer from rapid overfitting to the feedback signals generated by the frozen reward models, which is known as over-optimization. By employing adversarial feedback learning, the trainable reward model (acting as the discriminator) can swiftly adapt to the distribution shift of the diffusion model output, significantly mitigating the over-optimization phenomenon, which enables the reward model to provide effective guidance throughout the optimization process, (ii) Adversarial training expands the time step of feedback learning optimization: Including the strong adversarial targets in the training process forces high-noise timesteps to generate clearer images via the adversarial objective, which enables the reward model to perform well even under a few denoising steps. As presented in Fig.9(b), after disabling the adversarial loss and retaining the optimization step containing the entire denoising process, the reward model fails to provide effective guidance for samples under fewer denoising steps without the adversarial training object due to the high-level noise, which results in poor inference results. With these two benefits, adversarial feedback learning significantly improves the generation quality of samples in lower inference steps and finally achieves superior acceleration performance.
For more ablation study of UniFL, please refer to the Appendix.
5.4 Generalization Study
To further verify the generalization of UniFL, we performed downstream tasks including LoRA, ControlNet, and AnimateDiff. Specifically, we selected several popular styles of LoRAs, several types of ControlNet, and AnimateDiff modules [18] and inserted them into our models respectively to perform corresponding tasks. As shown in Fig.10 and Fig.8(b), our model demonstrates excellent capabilities in style learning, controllable generation, and video generation.
6 Discussion and Limitations

### v3

@@ 5.3 Ablation Study
To validate the effectiveness of our design, we systematically remove one component at a time and conduct a user study. The results are summarized in Fig.6 (a). In the subsequent sections, we will further analyze each component. More results are presented in the Appendix.
Figure 5: (a) Illustration of PeFL with instance segmentation model (SOLO). (b) Effect of PeFL on structure optimization.
Figure 6: (a) Design components ablation of UniFL. (b) Visualization of decoupled and non-decoupled aesthetic feedback learning results.
Superiority of PeFL. As depicted in Fig.5 (a), PeFL leverages the instance segmentation model to capture the overall structure of the generated object effectively. By identifying structural defects, such as the distorted limbs of the little girl, the broken elephant, and the missing skateboard, PeFL provides more precise feedback signals for diffusion models. Such fine-grained flaws can not be recognized well with ReFL due to its global and coarse preference feedback, instead, the exploited professional visual perception provides more detailed and targeted feedback. As presented in Fig.5 (b), the PeFL significantly boosts the object structure generation (e.g. the woman’s glasses, ballet dancer’s legs). It is also demonstrated by the notable performance drop (71.9% vs 28.1%) when disabling the PeFL.
Figure 7: Analysis on the $\alpha_{d}$. Left: reward scores distribution on 5k validation preference image pairs with our final chosen values highlighted. Right: ablation on the $\alpha_{d}$ on color and detail reward.
Multiple Aspects Optimization with PeFL. PeFL exploits various perceptual models to improve some particular visual aspects of the diffusion model and can easily be extended to multi-aspect optimization. As illustrated in Fig.8, the simultaneous incorporation of two distinct optimization objectives (style and structure optimization) does not compromise the effectiveness of each other. Take the prompt a baby Swan, graffiti as an example, integrating the style optimization via PeFL upon the base model successfully aligns the image with the target style. Further integrating the structure optimization objective preserves the intended style while enhancing the overall structural details (e.g. the feet of the Swan).
Figure 8: Incorporating the style and structure optimization objectives simultaneously with PeFL results in no effectiveness degeneration of each other.
Necessity of Decoupling Design. We conducted an experiment that finetuned the SD1.5 using the same prompt set but a global aesthetic reward model trained with all dimensions’ collected aesthetic preference data. As depicted in Fig.6 (b), the generated images are more harmonious and have an artistic atmosphere with the decoupled aesthetic reward tuning and are preferred by more 17% individuals than the non-decoupled counterpart. This can be attributed to the ease of abstract aesthetic learning with the decoupling design. Moreover, it also can be found that aesthetic feedback learning with actively selected prompts leads to a higher preference rate (54.6% vs 45.4%) compared with the random prompts. Further analysis of the prompt selection can be found in the Appendix.B.2.
Selection of Hinge Coefficient $\alpha_{d}$. We select the hinge coefficient for each aesthetic reward model based on their reward distributions on the validation set. As illustrated in Fig.7 (left). there are clear margins in the reward scores between preferred and unpreferred samples. Moreover, such margin varies across these dimensions, emphasizing the necessity of the decoupled design. Empirically, we set $\alpha_{d}$ to the average reward scores of the preferred samples to encourage the diffusion model to prioritize generating samples with higher reward scores. Fig.7 (right) demonstrates that setting a small hinge for the "color" reward resulted in only minor improvement, while substantial coefficients led to image oversaturation. Optimal results were achieved by selecting a coefficient close to the average reward score of the preferred samples. A similar trend was observed for layout and lighting aesthetics from our experiments, except for the "detail" dimension. Interestingly, a slightly lower coefficient sufficed for satisfactory detail optimization, as a higher coefficient introduced more background noise. This could be attributed to the significant reward score difference between preferred and unpreferred samples, where a high coefficient could excessively guide the model toward the target reward dimension.
Analysis on Adversarial Feedback Learning. We analyzed the mechanism for the acceleration behind our adversarial feedback learning and found that (i) Adversarial training enables the reward model to provide guidance continuously. As shown in Fig.9 (a), the diffusion model offer suffers rapid overfitting when frozen reward models, known as reward hacks. By employing adversarial feedback learning, the trainable reward model (acting as the discriminator) can swiftly adapt to the distribution shift of the diffusion model output, significantly mitigating the over-optimization phenomenon, and allowing the reward to provide effective guidance for a longer duration. (ii) Adversarial training expands the time step of feedback learning optimization. The adversarial objective poses a strong constraint to force high-noise timesteps to generate clearer images, which allows the samples across all denoising steps to be rewarded properly. As presented in Fig.9 (b), when disabling the adversarial objective while retaining the full optimization timesteps during rewarding, the reward model fails to provide effective guidance for samples under fewer denoising steps due to the high-level noise, which results in poor generation results. With these two benefits, adversarial feedback learning significantly improves the generation quality of samples in lower inference steps and achieves superior acceleration performance ultimately. Notably, as shown in Fig.6 (a), the image generated with 4-step inference retains similar visual quality with 20-step inference (52.5% vs 47.4%) after going through the second stage training, which demonstrates the superiority of UniFL in acceleration.
Figure 9: Analysis of the benefits of adversarial training. (a) It enables a longer optimization time for the reward model. (b) It enables the image under low denoising steps to be rewarded correctly. The red rectangle means incorporating the adversarial training objective.
Ablation on Acceleration Steps. We examine the acceleration capacity of UniFL under various inference steps, ranging from 1 to 8 as illustrated in Fig.10. Generally, UniFL performs exceptionally well with 2 to 8 inference steps with superior text-to-image alignment and higher aesthetic quality. The LCM method is prone to generate blurred images when using fewer inference steps and requires more steps (e.g., 8 steps) to produce satisfied images. However, both UniFL and LCM struggle to generate high-fidelity images with just 1-step inference, exhibiting a noticeable gap compared to SDXL-Turbo (e.g., the Labradoodle), which is intentionally designed and optimized for an extremely low-step inference regime. Therefore, there is still room for further exploration to enhance the acceleration capabilities of UniFL towards 1-step inference.
Figure 10: Ablation on different inference steps of UniFL.
6 Conclusion
We propose UniFL, a framework that enhances visual quality, aesthetic appeal, and inference efficiency for latent diffusion models from the unified feedback learning perspective. By incorporating perceptual, decoupled, and adversarial feedback learning, UniFL can be applied to various latent diffusion models, such as SD1.5 and SDXL, and exceeds existing methods in terms of both generation quality enhancement and inference acceleration.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (NO. 62322608), in part by the CAAI-MindSpore Open Fund, developed on OpenI Community, in part by the Open Project Program of State Key Laboratory of Virtual Reality Technology and Systems, Beihang University (No.VRLAB2023A01).
Appendix A Extend Details of Perceptual Feedback Learning
A.1 Additional Examples of PeFL
The proposed perceptual feedback learning (PeFL) is highly flexible, allowing it to utilize different existing visual perception models to offer targeted visual quality feedback on particular aspects. To showcase the scalability of PeFL, we present two additional case studies where PeFL is employed to optimize style and layout generation.
i) Style: To effectively capture image style and provide feedback on style generation, we utilize the VGG-16 [56] model to encode image features and extract visual style using the well-established gram matrix in style transfer. Furthermore, we have curated a substantial dataset of approximately 150,000 high-quality artist-style text images. We leverage this dataset to conduct PeFL for style optimization. The objective of the optimization can be formulated as follows:
$\mathcal{L}^{\mathrm{style}}_{\mathrm{pefl}}(\theta)=\mathbb{E}_{x_{0}\sim\mathcal{D},x^{\prime}_{0}\sim G(x_{t_{a}})}\|\mathrm{Gram}(V(x^{\prime}_{0}))-\mathrm{Gram}(V(x_{0}))\|_{2},$ (8)
where $V$ is the VGG network, and $\mathrm{Gram}$ is the calculation of the gram matrix. We validate the effect of PeFL in style optimization based on SD1.5 and SDXL. Note that due to the newly introduced artist-style dataset, we compare our method with the DMs fine-tuned with the same style dataset via pre-train loss to ensure a fair comparison. As depicted in Fig.11, the PeFL significantly boosts style generation (e.g. ’frescos’, ’impasto’ style), enabling the model to generate the image with a more aligned style compared with applied pre-train loss (MSE loss). We further conduct the quantitive experiment to evaluate the effectiveness of PeFL on style optimization. Specifically, we collect 90 prompts about style generation and generate 8 images for each prompt. Then, we manually statistic the rate of correctly responded generation to calculate the style response rate. As shown in Tab.2, it is clear that the style PeFL greatly boosts the style generation on both architectures thanks to the superior style feedback provided by the VGG extracted feature, especially for SD1.5 with about 15% improvement. In contrast, leveraging naive diffusion pre-train loss for fine-tuning with the same collected style dataset suffers limited improvement due to stylistic abstraction missing in latent space.
Figure 11: Style optimization of PeFL on SD1.5 and SDXL.
Model Style Response Rate
SD1.5 30.55%
SD1.5 + Style Pretrain 35.25%
SD1.5 + Style PeFL 45.14%
SDXL 66.67%
SDXL + Style Pretrain 68.34 %
SDXL + Style PeFL 75.27%
Table 2: Quantitive performance of PeFL in style generation.
ii) Layout: Generally, the semantic segmentation map characterizes the overall layout of the image as shown in Fig.12 (a). Therefore, semantic segmentation models can serve as a better layout feedback provider. Specifically, we utilize the visual semantic segmentation model to execute semantic segmentation on the denoised image $x^{\prime}_{0}$ to capture the current generated layout and supervise it with the ground truth segmentation mask and calculate semantic segmentation loss as the feedback on the layout generation:
$\mathcal{L}^{\mathrm{layout}}_{\mathrm{pefl}}(\theta)=\mathbb{E}_{x_{0}\sim\mathcal{D},x^{\prime}_{0}\sim G(x_{t_{a}})}\mathcal{L}_{\mathrm{semantic}}(m_{s}(x^{{}^{\prime}}_{0}),\mathrm{GT}(x_{0}))$ (9)
where $m_{s}$ represents the semantic segmentation model, $\mathrm{GT}(x_{0})$ is the ground truth semantic segmentation annotation and the $\mathcal{L}_{\mathrm{semantic}}$ is the semantic segmentation loss depending on the specific semantic segmentation model. We conducted an experiment on PeFL layout optimization based on SD1.5. Specifically, we utilize the COCO Stuff [57] with semantic segmentation annotation as the semantic layout dataset and DeepLab-V3 [58] as the semantic segmentation model. The results are presented in Fig.12 (b). It demonstrates that the PeFL significantly improves the layout of the generated image, for instance, the bear on the bed in a diagonal layout. Note that here we focus on the objective layout generation that is explicitly mentioned in the prompts, for example, ‘stands‘, ‘overlooking‘, and ‘sit at‘. As a comparison, the layout reward model used in aesthetic feedback learning primarily emphasizes the subjective composition from the aesthetic angle, which may not be described clearly by the textual prompt. We further conduct the user study to evaluate the effectiveness of PeFL with the semantic segmentation model quantitatively. As shown in Fig.13 (b), we are surprised to find that the image details also observed a significant boost in addition to the improvement in the layout generation. This probably stems from the dense per-pixel feedback from the semantic segmentation objective.
Indeed, PeFL is an incredibly versatile framework that can exploit a wide range of visual perceptual models, such as OCR models [59, 60] and edge detection models [61, 62], to boost the performance of LDMs. Furthermore, we are actively delving into utilizing the visual foundation model, such as SAM [63], which holds promising potential in addressing various visual limitations observed in current diffusion models.
Figure 13: (a) The user study results on the ablation of different instance segmentation models in PeFL during structure optimization. PeFL (SOLO): PeFL fine-tune SD1.5 with SOLO as the instance segmentation model. PeFL (Mask2Former): PeFL fine-tune SD1.5 with Mask2Former as the instance segmentation model. (b) The user study results on the effect of PeFL layout optimization. Dark Green: SD1.5 with PeFL, Light Green: SD1.5 without PeFL.
A.2 Ablation on Visual Perceptual Model
PeFL utilizes various visual perceptual models to provide visual feedback in specific dimensions to improve the visual generation quality on particular aspects. Different visual perceptual models of a certain dimension may have different impacts on the performance of PeFL. Taking the structure optimization of PeFL as an example, we investigated the impact of the accuracy of instance segmentation models on PeFL performance. Naturally, the higher the precision of the instance segmentation, the better the performance of structure optimization. To this end, we choose the Mask2Former [64], another representative instance segmentation model with state-of-the-art performance to achieve structure optimization with PeFL. The results are shown in Fig.16 (a) and Fig.13 (a). It is intriguing to note that the utilization of a higher precision instance segmentation model does not yield significantly improved results in terms of performance. We speculate it lies in the different architectures of the instance segmentation of these two models. In SOLO [50], the instance segmentation is formulated as a pixel-wise classification, where each pixel will be responsible for a particular instance or the background. Such dense supervision fashion enables the feedback signal to better cover the whole image during generation. In contrast, Mask2Former [64] takes the query-based instance segmentation paradigm, where only a sparse query is used to aggregate the instance-related feature and execute segmentation. This sparse nature of the query-based method makes the feedback insufficient and leads to inferior fine-tuning results. We leave further exploration of how to choose the most appropriate visual perceptual model for feedback tuning to future work.
A.3 Generalization of PeFL with Close-set Perceptual models
Figure 14: Generalization of PeFL with SOLO. The generation of the concepts not included in COCO (e.g. ostrich, robot, cones) is also improved after PeFL optimization.
We utilize the SOLO instance segmentation model trained with close-set dataset COCO for PeFL structure optimization. One may be concerned that this will lead to limited concepts that PeFL can optimize (i.e. only the COCO concept). However, on the one hand, although we apply the SOLO instance segmentation model trained on the COCO dataset (80 categories), we observe that PeFL exhibits exceptional generalization capability and the generation performance of many concepts not shown in the COCO dataset is also boosted significantly as shown in Fig.14. We believe LDM can be guided to learn general and reasonable structure generation via PeFL optimization. On the other hand, the proposed perceptual feedback learning is a very flexible framework, and it is very straightforward to replace the close-set model SOLO with other open-set instance segmentation models such as Ground-SAM to achieve further improvement for the concepts in the wild(e.g. concepts in LAION dataset).
Appendix B Extend Details of Decoupled Feedback Learning
B.1 Aesthetic Preference Data Collection
Figure 15: Decoupled aesthetic feedback data examples. The preferred samples are highlighted with red rectangles.
We break down the general and coarse aesthetic concept into more specific dimensions including color, layout, detail, and lighting to ease the challenge of aesthetic fine-tuning. We then collect the human preference dataset along each dimension. Specifically, we employ the SDXL [40] as the base model and utilize the prompts from the MidJourney [3, 65] as input, generating two images for each prompt. Subsequently, we enlist the expertise of 4 to 5 professional annotators to assess and determine the superior image among the generated pair. Given the inherently subjective nature of the judgment process, we have adopted a voting approach to ascertain the final preference results for each prompt. Finally, we curate 30,000, 32,000, 30,000, and 30,000 data pairs for the color, layout, detail, and lighting dimensions, respectively. Examples of the collected aesthetic feedback data of different dimensions are visually presented in Fig.15.
B.2 Active Prompt Selection
Prompt Selection Process. We introduce an active prompt selection strategy designed to choose the most informative and diverse prompts from a vast prompt database. The comprehensive implementation of this selection strategy is outlined in the Algorithm.2. Our strategy’s primary objective is to select prompts that offer maximum information and diversity. To accomplish this, we have devised two key components: the Semantic-based Prompt Filter and the Nearest Neighbor Prompt Compression. The semantic-based prompt filter is designed to assess the semantic relationship embedded within the prompts and eliminate prompts that lack substantial information. To accomplish this, we utilize an existing scene graph parser [^1] https://github.com/vacancy/SceneGraphParser as a tool to parse the grammatical components, such as the subjective and objective elements.
Algorithm 2 Active prompt selection for decoupled aesthetic feedback learning
0:  Initial prompt database: $\mathcal{D}$; number of desired prompts: $N$; $\tau_{1}$ and $\tau_{2}$: relation and similarity threshold.
0:  The selected prompt set: $\mathcal{SP}$
1:  $\mathcal{P}$ = $\varnothing$
2:  # Semantic-based Prompt Filter
3:  for  $p_{i}\in\mathcal{D}$ do
4:     $\mathcal{SR}$leftarrow SemanticParser($p_{i}$)
5:     if $|\mathcal{SR}|>\tau_{1}$ then
6:        $\mathcal{P}\leftarrow p_{i}$ // choose informative prompt
7:     end if
8:  end for
9:  # Nearest Neighbor Prompt Compression
10:  $I$leftarrow shuffle(range(len(| $\mathcal{P}$|)))
11:  $R$leftarrow False // set the removed prompt array
12:  $S$leftarrow $\varnothing$ // set the selected prompt index
13:  Dist, Inds $\leftarrow$ KNN($R$, $k$) // K-nearest neighbor of each prompt
14:  for index $I_{i}\in I$ do
15:     if not $R[I_{i}]$ and $I_{i}$ not in $S$  then
16:        $\mathcal{S}\leftarrow I_{i}$ // append the selected prompt
17:        dist, inds = Dists[$I_{i}$], Inds[$I_{i}$] // K-nearest neighbor similarities
18:        for index $d_{i}$in inds do
19:           if dist[$d_{i}$] > $\tau_{2}$ then
20:              $R[d_{i}]$ = True
21:           end if
22:        end for
23:     end if
24:  end for
25:  $\mathcal{SP}$leftarrow RandomSelect($\mathcal{P}$, $S$, $N$) // randomly select $N$ diverse prompts according the retained index
26:  return  $\mathcal{SP}$
The scene graph parser also generates various relationships associated with the subjective and objective, including attributes and actions. We then calculate the number of relationships for each subjective and objective and select the maximum number of relationships as the measurement of the information amount encoded in the prompt. A higher number of relationships indicates that the prompt contains more information. We filter out prompts that have fewer than $\tau_{1}=1$ relationships, which discard the meaningless prompt like ‘ff 0 0 0 0‘ to reduce the noise of the prompt set. Upon completing the filtration process, our next objective is to select a predetermined number of prompts that exhibit maximum diversity. To achieve this, we adopt an iterative process to achieve this objective. In each iteration, we randomly select a seed prompt and subsequently suppress its nearest neighbor [^2] https://github.com/facebookresearch/faiss prompts that have a similarity greater than $\tau_{2}=0.8$ as illustrated in Fig.16 (b). The next iteration commences with the remaining prompts, and we repeat this process until the similarity between the nearest neighbors of all prompts falls below the threshold $\tau_{2}$. Finally, we randomly select the prompts, adhering to the fixed number required for preference fine-tuning.
Analysis of the Actively Selected Prompts. As illustrated in Fig.6 (a) in the main paper, our strategically chosen prompts yield superior performance in aesthetic feedback learning. To further comprehend the advantage of this design, we present the training loss curve in Fig.16 (c), comparing the use of actively selected prompts versus random prompts. It clearly shows that the diffusion model rapidly overfits the guidance provided by the reward model when using the randomly selected prompts, ultimately resulting in the loss of effectiveness of the reward model quickly. One contributing factor to this phenomenon is the distribution of prompts for optimization. If the prompts are too closely distributed, the reward model is forced to frequently provide reward signals on similar data points, leading to the diffusion model rapidly overfitting and collapsing within a limited number of optimization steps. We statistic the average nearest embedding similarity (ANS) of the 100K prompts randomly selected from DiffusionDB [48] by calculating the cosine similarity between each prompt with its most similar prompt within the embedding space and taking the average over all the prompts. The ANS of the randomly selected prompts is approximately 0.89, which delivery highly redundant. As a comparison, the prompts selected by our strategy exhibit considerable diversity with ANS of 0.73, enabling a more balanced and broad reward calculation, which eases the over-fit significantly. Therefore, with the actively selected prompts, the diffusion model obtains a more comprehensive feedback signal and can be optimized toward human preference more efficiently.
Figure 16: (a) The visual comparison between the PeFL structure optimization with different instance segmentation models. (b) Illustration of Nearest Neighbor Prompt Compression. (c) Training loss curve when utilizing different prompts for decoupled aesthetic feedback learning.
Appendix C Generalization Study
To further verify the generalization of UniFL, we performed downstream tasks including LoRA, and ControlNet Specifically, we selected several popular styles of LoRAs [7], and several types of ControlNet [9] and inserted them into our models respectively to perform corresponding tasks. As shown in Fig.17, our model demonstrates excellent capabilities in style adaptation and controllable generation.
Appendix D More Visualization Results
@@
Figure 18: More visual comparison with different methods. UniFL yields better prompt-aligned and visually appealing results than other methods. Our results are highlighted with red rectangles.
Appendix E Discussion and Limitations